I have one gallon of acid 30% solution . I need to reduce it to 5% acid. How much water should I add?
Answer by jim_thompson5910(35256)

"I have one gallon of acid 30% solution"

So you have 1*0.30 = 0.30 gallons of pure acid.

-------------------------------------------------------

This means

(amount of pure acid)/(total amount of solution) = (0.30)/(1) = 0.30

which points back to the 30% acid solution

-------------------------------------------------------

Let x = amount of water to add (amount in gallons)

If you add x gallons of water to the solution, you get 1+x total gallons (acid + water)

so we now have

(amount of pure acid)/(total amount of solution) = (0.30)/(1+x)

You then take (0.30)/(1+x) and set that equal to 0.05 because you want to force the final solution to be 5% acid.

So,

(0.30)/(1+x) = 0.05

Now solve for x

(0.30)/(1+x) = 0.05
0.30 = 0.05(1+x)
0.30 = 0.05+0.05x
0.30-0.05 = 0.05x
0.25 = 0.05x
0.25/0.05 = x
5 = x
x = 5


You need to add 5 gallons of pure water to dilute the 30% acid solution to a 5% acid solution.